Good day tutor
Please I need your assistance on these questions below...god bless
1. Subtract 3×-8×2+17 from
2×2+7×-10
2.divide 3p3-7p2+10p-48 by p3
3.5×3x4×÷10×4
4. Calculate the value of r in the formula in (b) if y =13 and x=17
5.evaluate without using a calculator 3sin 30°+tan 45°-2cos 60°

Answer by rothauserc(4718)

1) (2x^2 +7x +10) - ((-8x^2) +3x + 17) =
2x^2 +7x +10 +8x^2 -3x -17 =
10x^2 +4x -7
**********************************************************************
2) (3p^3 -7p^2 +10p -48) / p^3 =
(3p^3/p^3) -(7p^2/p^3) +(10p/p^3) -(48/p^3) =
(3) -(7/p) +(10/p^2) -(48/p^3) =
-48/p^3 +10/p^2 -7/p +3
**********************************************************************
3) (5x * 3x *4x) / (10 * x^4) =
(60 * x^3) / (10 * x^4) =
6 / x
*********************************************************************
4) where is formula in (b)?
*********************************************************************
5) (3 * sin(30)) + tan(45) -(2 * cos(60)) =
from the unit circle we know sin 30 = 1/2, cos 60 = 1/2 and tan 45 = 1
(3 * (1/2) + 1 -(2 * (1/2) =
3/2
*********************************************************************